Title: 2 Blazars
Session: Feb 1997 (session 1)
Frequency: 18 cm
Obs time (UT): EU 18:30(27/02)-08:30(28/02)
EU Mk3 Tapes/St.: 3
To give feedback on this experiment, click here
Medicina (E.Arbizzani at Fri Feb 28 03:45:43 EST 1997)Success
Scan 7:40-7:53 some tsys problems
Westerbork (Tony Foley at Fri Feb 28 09:00:58 EST 1997) Minor failures
Head stuk on pass 12, 0017-0030UT.
MAJOR problem with first synthesizer set wrongly from 0254UT to the
end (VLBI in conflict with online system). Frequency will be
2.7MHz too high.
Onsala (Antonis Polatidis at Fri Feb 28 11:29:02 EST 1997) Minor failures
Between 1830-1843 and 1908-1920 we recorded nor,nor instead for nor,alt so even VCs contain the wrong data. Apologies for the error. Lost first scan of the third tape due to loss of vacuum. No other known problems
Shanghai (Huang Xinyong at Wed Mar 5 05:14:34 EST 1997) Success
Clock offset = 0.51 - 0.53 msec. No problems, weather overcast.
Jodrell (Alastair Gunn at Tue Mar 18 12:27:47 MET 1997) Minor failures
POLARISATION ARE CROSSED therefore no useful data.
Calibration (Fredrik T Rantakyro at Thu May 8 11:37:33 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the
Feb97 directory with the name em24.antab. This file contains the
preliminary calibration results and may be updated when I receive
updated information from the telescopes and the results from the
datareduction of the calibration experiments.
MPIfR (Walter Alef at Wed Jan 7 12:09:31 MET 1998) Failure
Crimea, Cambridge and Noto did not observe. No fringes with Urumqi. Wrong pol.
Jodrell and Onsala. Dual pol Torun. 30% cross talk Westerbork. More than 50% of